# {textblock*}

### v1

@@ 1 Introduction
In contrast, a more recent and practical (cf. Section 2) view considers probabilistic robustness (PR) [43, 44, 42, 34, 11, 3, 40, 46, 41, 48]—it answers:
@@ 2 Definition of Probabilistic Robustness
It is evident that cases (a), (b), and (c) in Fig. 1 focus on extreme scenarios, where robustness is evaluated based on the existence of a deterministic AE, reflecting a worst-case perspective. This approach, originally termed “adversarial” robustness, is particularly relevant in malicious contexts and security studies. While a comprehensive review of AR is out of the scope of this paper, we refer readers to surveys [22, 31, 53, 9]. In contrast, case (d) adopts a probabilistic view, assessing overall local robustness by considering the presence of AEs statistically. Compared to AR, this probabilistic view of PR is more practical for real-world applications, as worst-case evaluations are often unnecessary, especially for large models operating in benign environments. Quantifying the proportion/risk of AEs is more relevant for reliability and safety, where ensuring the AE risk remains below an acceptable threshold is sufficient, rather than requiring it to be exactly zero [43, 42, 46, 48].
We note that, although some studies generate perturbations stochastically, this does not inherently imply a focus on PR. For example, randomised smoothing techniques [10] aim to certify a safe perturbation radius (i.e., case (b) in Fig. 1) by replacing standard model inference with the averaging of predictions over a set of randomly perturbed inputs.
@@
More studies on bridging local PR to global properties are presented in Sec. 5.
@@ 4 Enhancement of Probabilistic Robustness
Despite the relatively greater focus on evaluating PR, to the best of our knowledge, the only work explicitly aimed at improving the PR of DL models is [42], which employs data augmentation. In this training scheme, [42] demonstrates—both empirically and theoretically—that simply adding perturbed inputs significantly enhances PR and leads to better generalisation (i.e., the robustness improvement in training set can be better generalised to the test set). However, this approach provides little to no improvement in AR. We believe that, akin to the well-established AT framework for AR, a dedicated AT framework for PR is necessary. While [42] incorporates data augmentation to enhance PR, a systematic AT approach for PR remains absent in the literature. Developing such an AT for PR scheme would enable the joint improvement of both PR and AR.
AT [19, 28, 2] is a widely used empirical approach for improving robustness, typically formulated as a “min-max” optimisation:
Definition 4 (AT for AR)
For a DL model $f_{\theta}$, trained on a dataset $\mathcal{S}$, consisting of pairs $(x,y)$ where $x\in\mathcal{X}$ is the input and $y\in\mathcal{Y}$ denotes the label, with loss function $\mathcal{L}(x,y;\theta)$, the objective of AT-AR can be formulated as:
$\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{S}}\left[\max_{\|\delta\|\leq\gamma}\mathcal{L}(x+\delta,y;\theta)\right],$ (4)
where $\delta$ represents the perturbation in a norm ball with radius $\gamma$ applied to $x$.
The “inner max” often empirically solved by adversarial attacks, e.g., PGD [28] has become a critical benchmark and is widely regarded as the standard method for AT in practice [2]. The objective of AT for AR is to identify the AEs that maximises model loss and incorporate it into AT, cf. top-middle illustration in Fig. 2. However, to enhance local PR, the goal shifts to identifying the widest peak—the region with the highest proportion of AEs in the loss landscape—rather than the tallest peak (cf. top-right illustration in Fig. 2). Incorporating the AE from this region into AT can potentially reduce loss for AEs in its vicinity, thereby lowering the overall AE proportion (cf. PR interpretation in Definition 1). In some cases, AT-WCR and AT-PR may select the same AE when the widest and tallest peaks coincide, as depicted in the second row of Fig. 2. The two scenarios illustrated in Fig. 2 representatively cover all possible cases, though additional peaks may exist in practice.
@@
The intuition behind the new “inner max” is to find an optimal perturbation $\delta$ that produces an AE $x^{\prime}=x+\delta$ that maximises the radius $k$ of a smaller norm-ball centred on this AE $x^{\prime}$, such that the PR within this $k$-sized norm-ball is 0, i.e., all inputs within this new and widest norm-ball are AEs. The high non-linearity and non-convexity of DL models makes it intractable to analytically solve both the established min-max in Eq. (4) and our new formulation in Eq. (5). Novel and efficient numerical algorithms to solve the new framed min-max problem of AT-PR will be an important research direction.
@@ 6 Challenges and Open Problems
We identify the following key challenges and open problems for future research.
@@
PR research has predominantly focused on image classification tasks, as highlighted in Table 1, with limited exploration into other domains. Only one study has investigated text-to-image models [48], and another has addressed XAI tasks. Expanding PR research to multi-modal and generative AI tasks is essential for achieving broader applicability and impact. However, unlike classification, which typically involve continuous image inputs and discrete categorical outputs, these tasks often deal with diverse modalities, e.g., images, text, and audio, as both inputs and outputs. A key open problem is the development of task-specific PR formulations that are computationally feasible while effectively capturing perturbations relevant to each modality.
@@
Unlike AR, which has well-established AT techniques formulated as a min-max optimisation in Eq. (4) , no dedicated AT scheme exists for PR. In this paper Sec. 4, we have reformulated the min-max optimisation specifically for PR, establishing a new theoretical foundation for PR-focused AT. Designing AT schemes with efficient numerical algorithms to solve this new optimisation problem is challenging and forms an important future work.
@@
A critical area for future research is the theoretical analysis and generalization bounds for PR optimised models. Understanding how PR generalises across diverse data distributions and model architectures remains an open challenge. This necessitates advancements in learning theories, particularly in areas such as Rademacher complexity and PAC-Bayesian bounds, to provide rigorous theoretical foundations.

### v2

@@ 1 Introduction
In contrast, a more recent and practical (cf. Section 2) view considers probabilistic robustness (PR) [46, 47, 45, 35, 11, 3, 43, 49, 39, 44, 52]—it answers:
@@ 2 Definition of Probabilistic Robustness
While PR in Eq. (1) represents the expected probability of detecting AEs, alternative tail-sensitive local PR measures can be defined, e.g., $\rho$-essential supremum ($\rho$-ess sup) for a given PR threshold and Conditional Value-at-Risk (CVaR) for tail averages [39], and Entropic Value-at-Risk (EVaR) for exponential tail bounds [50].
It is evident that cases (a), (b), and (c) in Fig. 1 focus on extreme scenarios, where robustness is evaluated based on the existence of a deterministic AE, reflecting a worst-case perspective. This approach, originally termed “adversarial” robustness, is particularly relevant in malicious contexts and security studies. While a comprehensive review of AR is out of the scope of this paper, we refer readers to surveys [23, 32, 57, 9]. In contrast, case (d) adopts a probabilistic view, assessing overall local robustness by considering the presence of AEs statistically. Compared to AR, this probabilistic view of PR is more practical for real-world applications, as worst-case evaluations are often unnecessary, especially for large models operating in benign environments. Quantifying the proportion/risk of AEs is more relevant for reliability and safety, where ensuring the AE risk remains below an acceptable threshold is sufficient, rather than requiring it to be exactly zero [46, 45, 49, 52, 39].
Although some studies generate perturbations stochastically, this does not imply a focus on PR, e.g., randomised smoothing [10] aim to certify a safe perturbation radius (i.e., case (b) in Fig. 1) by replacing standard model inference with the averaging of predictions over a set of randomly perturbed inputs.
@@
Again, the “inner-expectation” can be replaced by other risk measures like $\rho$-ess sup and CVaR in [39], and EVaR in [50].
@@ 3 Evaluation of Probabilistic Robustness
ICML’22 [39] $\rho$-ess sup CVaR $L_{p}$ noise stochastic gradient decent, simple Monte Carlo Img. classific.
@@
IJCAI’24 [50] EVaR $L_{p}$ noise adaptive stochastic search, simple Monte Carlo Img. classific.
@@ 4 Enhancement of Probabilistic Robustness
Despite the relatively greater focus on evaluating PR, to the best of our knowledge, limited works explicitly aimed at improving the PR of DL models are [45, 39, 50]. [45] demonstrates—both empirically and theoretically—that simply adding a randomly perturbed inputs significantly enhances PR and leads to better generalisation (i.e., the robustness improvement in training set can be better generalised to the test set). Both [39, 50] design loss functions using common risk measures—$\rho$-essential supremum, Conditional Value-at-Risk, and Entropic Value-at-Risk—then train models to minimize them on the training dataset.
However, these three approaches provide quite limited improvement in AR. We believe that, akin to the well-established AT framework for AR, a dedicated AT framework for PR is necessary, which remains absent in the literature. Developing such an AT for PR scheme would enable significant joint improvement of both PR and AR. AT [20, 29, 2] is a widely used empirical approach for improving robustness, typically formulated as a “min-max” optimisation. The “inner max” often empirically solved by adversarial attacks, e.g., PGD [29] has become a critical benchmark and is widely regarded as the standard method for AT in practice [2]. The objective of AT for AR is to identify the AEs that maximises model loss and incorporate it into AT, cf. top-middle illustration in Fig. 2. However, to enhance local PR, the goal shifts to identifying the widest peak—the region with the highest proportion of AEs in the loss landscape—rather than the tallest peak (cf. top-right illustration in Fig. 2). Incorporating the AE from this region into AT can potentially reduce loss for AEs in its vicinity, thereby lowering the overall AE proportion (cf. PR interpretation in Definition 1). In some cases, AT-WCR and AT-PR may select the same AE when the widest and tallest peaks coincide, as depicted in the second row of Fig. 2. The two scenarios illustrated in Fig. 2 representatively cover all possible cases, though additional peaks may exist in practice.
@@
The intuition behind the new “inner max” is to find an optimal perturbation $\delta$ that produces an AE $x^{\prime}=x+\delta$ that maximises the radius $k$ of a smaller norm-ball centred on this AE $x^{\prime}$, such that the PR within this $k$-sized norm-ball is 0, i.e., all inputs within this new and widest norm-ball are AEs. The high non-linearity and non-convexity of DL models makes it intractable to analytically solve both the established min-max for AR and our new formulation in Eq. (4). Novel and efficient numerical algorithms to solve the new framed min-max problem of AT-PR will be an important research direction.
@@ 6 Challenges and Open Problems
PR research has predominantly focused on image classification tasks, as highlighted in Table 1, with limited exploration into other domains. Only one study has investigated text-to-image models [52], and another has addressed XAI [22]. Expanding PR research to multi-modal and generative AI tasks is essential for achieving broader applicability and impact. However, unlike classification, which typically involve continuous image inputs and discrete categorical outputs, these tasks often deal with diverse modalities, e.g., images, text, and audio, as both inputs and outputs. A key open problem is the development of task-specific PR formulations that are computationally feasible while effectively capturing perturbations relevant to each modality. Furthermore, a recent work [19] introduce PR for conformal prediction techniques.
@@
Unlike AR, which has well-established AT techniques, no dedicated AT scheme exists for PR. In this paper Sec. 4, we have reformulated the min-max optimisation specifically for PR, establishing a new theoretical foundation for PR-focused AT. Designing AT schemes with efficient numerical algorithms to solve this new optimisation problem is challenging and forms an important future work.
@@
A critical area for future research is the theoretical analysis for PR optimised models. Understanding how PR generalises across diverse data distributions and model architectures remains an open challenge. This necessitates advancements in learning theories, particularly in areas such as Rademacher complexity and PAC-Bayesian bounds [37], to provide rigorous theoretical foundations.